10 grams of metal oxide contains 6 grams of Oxygen.the equivalent weight of  the metal carbonate is
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
let Metal oxide be M2Ox .
then Total moles of compound = 10/(2M + 16x)
moles of M = > 20/(2M+16x) = 4/2M
moles of O => 10x/(2M + 16x) = 6/16x
solving both eqns . we get,
M = 5.09. and x= 1 approx.
Hence eq. weight for metal carbonate => M2CO3
=> 2M + 48 + 12
=> 10.18 +  48 + 12
=> 70.18.